OpenStudy (anonymous):

I'm supposed to be able to solve ln1 without using a calculator. ...how?

hartnn (hartnn):

if \(\ln a = b, then \: a=e^b \) let ln 1 =x then e^x = 1 which value of 'x' you think, satisfies this ?

OpenStudy (anonymous):

... 0 :)

hartnn (hartnn):

thas correct, infact in any base of log, the answer will be 0 ln 1=0 , log1=0 , log_n 1=0 because anything^x = 1 gives you x=0